OpenStudy (anonymous):

Express 3x+4y=12 in slope-intercept form.

OpenStudy (anonymous):

y = 3/4x - 3

OpenStudy (anonymous):

@Bkfrostyblue is there a way to show any work for this problem?

OpenStudy (anonymous):

yah 3x - 4y = 12 subtract 3x from both sides -4y = -3x + 12 -4y/4 = -3x/4 + 12/4 (divide everything by 4) -y = -3/4x + 3 (multiply -1 to everything just to make it look simpler) y = 3/4x - 3

OpenStudy (anonymous):

Ok thank you Very Much. @Bkfrostyblue

OpenStudy (anonymous):

no problem ;)